# Self-consistent electrostatic formalism of bulk electrolytes based on the asymmetric treatment of the short- and long-range ion interactions

### v1

@@ Abstract
ABSTRACT: We predict the thermodynamic behavior of bulk electrolytes from an ionic Hard-Core (HC) size-augmented self-consistent formalism incorporating asymmetrically the short- and long-range ion interactions via their virial and cumulant treatment, respectively. The characteristic splitting length separating these two ranges is obtained from a variational equation solved together with the Schwinger-Dyson (SD) equations. Via comparison with simulation results from the literature, we show that the asymmetric treatment of the distinct interaction ranges significantly extends the validity regime of our previously developed purely cumulant-level Debye-Hückel (DH) theory. Namely, for monovalent solutions with typical ion sizes, the present formalism can accurately predict mutually the HC-dominated liquid pressure and the electrostatic correlation-driven internal energies up to molar concentrations. We evaluate as well the screening length of the liquid and investigate the deviations of the intermolecular interaction range from the DH length. We show that in fair agreement with simulations and experiments, our theory can also reproduce the overscreening and underscreening effects occurring respectively in mono- and multivalent electrolytes.
@@ I Introduction
In order to extend the electrostatic coupling regime covered by the CCDH formalism, in this article, we develop a self-consistent DH (SCDH-I) theory of bulk electrolytes based on the asymmetric treatment of the short- and long-range electrostatic fluctuations. In Sec. II, we derive the field-theoretic partition function of the liquid embodying the corresponding splitting together with the pairwise HC interactions, and obtain a new variational equation solved by the splitting parameter separating these distinct interaction ranges. Then, we derive the SD equations relating this variational identity to the physical parameters of the system. Finally, we calculate the total correlation function required for the computation of the thermodynamic functions and the screening length of the electrolyte. The thermodynamic averages involved in these formally exact identities are evaluated via an hybrid calculation scheme that consists in incorporating asymmetrically the electrostatic fluctuations of short- and long-wavelengths within the virial and cumulant approximations, respectively. This scheme corresponding to a variational upgrade of our electrostatic WC-level CCDH formalism Buyuk2024 allows to avoid the cumulant-level WC treatment of the short-range ion interactions.
In Sec. III, we compare our predictions with hypernetted-chain (HNC) and MC data for various ion sizes and valencies. We find that the asymmetric variational treatment of the distinct interaction modes significantly improves the accuracy of the CCDH formalism in predicting mutually the HC-dominated osmotic pressure and the electrostatic correlation-driven internal energies. This upgrade is the main progress of our work. Via additional comparison with simulations and experiments, we equally show that our approach can also reproduce with reasonable accuracy the underscreening and overscreening effects occurring in mono- and multivalent solutions.
@@ II.6.1 Cumulant-level calculation of the ion fugacities
$-t\Lambda_{i}q_{i}^{2}\int\frac{\mathrm{d}^{3}\mathbf{r}_{1}\mathrm{d}^{3}\mathbf{r}_{2}}{2}\left[G_{{\rm l},0}^{-1}-v_{\rm c}^{-1}\right]_{\mathbf{r}_{1},\mathbf{r}_{2}}$
@@
with the rescaled fugacity $\Lambda_{i}=\lambda_{i}\;e^{-q_{i}^{2}\delta G_{{\rm l},0}(0)/2}$, the Mayer function
@@
$+tq_{i}^{2}n_{i}\int\frac{\mathrm{d}^{3}\mathbf{r}_{1}\mathrm{d}^{3}\mathbf{r}_{2}}{2}\left[G_{{\rm l},0}^{-1}-v_{\rm c}^{-1}\right]_{\mathbf{r}_{1},\mathbf{r}_{2}}$
@@ II.6.3 Cumulant-level evaluation of the correlation functions
Figure 1: (Color online) (a)-(c) Internal energy (83) and (b)-(d) pressure (84) of 1:1 solutions against salt concentration at two different HC sizes indicated in the legends. The temperature and the dielectric constant of the liquid are $T=300$ K and $\varepsilon_{\rm w}=78.5$. Solid curves: SCDH-I theory. Dashed curves: CCDH formalism of Ref. Buyuk2024 . Dotted curves: DHLLs. Red squares: HNC data from Figs. 3 and 4 of Ref. AttardPRE .
@@ III Results and discussion
In this section, we compare the predictions of the present formalism with numerical and experimental data from the literature. For the interpretation of the results reported herein, it is essential to note that the CCDH theory of Ref. Buyuk2024 follows from the present formalism in the limit of vanishing splitting parameter $\sigma\to 0$ where the short-range potential (18) in Eq. (II.6.2) vanishes ($v_{\rm s}(r)\to 0$), and the Gaussian component of the long-range correlator reduces to the electrostatic WC-level DH potential, i.e. $G_{{\rm l},0}(r)\to\ell_{\rm B}e^{-\kappa_{0}r}/r$, where $\kappa_{0}=\sqrt{8\pi\ell_{\rm B}n_{i}}$ is the DH screening parameter. Thus, the CCDH theory corresponds to the electrostatic WC limit of the SCDH-I approach. This implies that the pure HC interactions are embodied by the SCDH-I formalism at the same level of accuracy as the CCDH theory.
@@
Figure 2: (Color online) Internal energy (83) against salt concentration. The ion valencies and sizes are indicated in the legends. The temperature and the dielectric constant are $T=298$ K and $\varepsilon_{\rm w}=78.5$. The curves have the same signification as in Fig. 1. The MC data (red squares) are from (a) Table I, (b) III, (c) IV, and (d) II of Ref. Valleau .
III.1 Excess energy and pressure
Figure 3: (Color online) (a)-(e) Internal energy (83) and (f)-(j) pressure (84) of monovalent ions against their packing fraction $\eta=\pi n_{i}d^{3}/3$ (lower axis) and concentration (upper axis) at the temperature $T=298$ K, the dielectric permittivity $\varepsilon_{\rm w}=78.5$, and the HC sizes indicated in the legends rem1 . MC data are from (a)-(e) Table I and (f)-(j) Table II of Ref. NetzMC .
@@ III.1 Excess energy and pressure
The excess energy density and pressure of a general electrolyte mixture can be expressed in terms of the total correlation function (80) as Buyuk2024
@@
Figs. 1(a)-(d) compare the excess energy (83) and pressure (84) of the SCDH-I approach (solid curves) with HNC data (red squares) for 1:1 liquids at two different ion sizes. The plots also display the CCDH predictions (dashed curves), and the DH limiting laws (DHLLs),
$\beta E_{{\rm ex,DH}}=-\frac{\kappa_{0}^{3}}{8\pi};\hskip 14.22636pt\beta P_{\rm DH}=\sum_{i=1}^{p}n_{i}-\frac{\kappa_{0}^{3}}{24\pi}$ (85)
@@
(dotted curves). The figures show that the variational scheme at the basis of the SCDH-I formalism significantly improves the agreement of the CCDH theory with HNC data. Namely, while the CCDH formalism extends the validity of the DHLLs from $n_{i}\sim 10^{-2}$ M to $n_{i}\sim 0.2$ M, the SCDH-I results agree quantitatively with HNC data up to $n_{i}\sim 1.0$ M. Hence, for monovalent ions with typical hydration radii $d\sim 4$, the asymmetric treatment of the short- and long-range ion interactions enables the SCDH-I approach to accurately predict both the charge correlation-driven internal energy and the HC-dominated liquid pressure up to the molar concentration regime.
@@
Fig. 2(a) confronts the theoretical internal energy predictions for 1:1 solutions with MC simulations over a larger concentration regime. The quantitative accuracy of the SCDH-I approach up to $n_{i}\sim 1.0$ M is equally confirmed by this plot. At larger concentrations $n_{i}\gtrsim 2.0$ M where the SCDH-I prediction overestimates the MC data, the qualitative agreement of the present formalism with simulations is significantly better than that of the CCDH approach exhibiting a trend reversal, and an unphysical rise of the energy previously reported in Ref. Buyuk2024 .
@@
Figs. 2(b)-(d) show that for multivalent electrolytes characterized by stronger electrostatic coupling, the quantitative agreement of the SCDH-I formalism with MC data is limited to submolar concentrations. However, in the vicinity of the molar concentration regime, the SCDH-I approach can qualitatively reproduce the trend and magnitude of the simulation data with a considerably better accuracy than the CCDH formalism. We finally note that in Figs. 2(d) for 2:2 liquids, the SCDH-I curve has been plotted up to $n_{i}\sim 3.0$ M where the numerical scheme for the solution of the variational Eq. (79) fails.
@@
We investigate now the effect of HC size on the accuracy of the SCDH-I formalism. In Fig. 3(a), the comparison with MC data shows that at the particularly large solute size $d=14.29\;{\rm{\AA}}>\ell_{\rm B}\approx 7.1$ where the strong attenuation of opposite charge attraction by HC repulsion gives rise to weak electrostatic coupling conditions, the SCDH-I formalism brings a moderate correction to the CCDH formalism. Then, in the typical case of smaller ion sizes $d<\ell_{\rm B}$ of Figs. 3(b)-(e) where the shorter interionic approach distances enhance the strength of the electrostatic correlations, the SCDH-I formalism substantially improves the quantitative accuracy of the CCDH approach, extending its validity limit by several factors into the molar concentration regime.
@@
In Figs. 3(f)-(j), we illustrate the effect of solute size on the liquid pressure (84). At the ion size $d=14.29$ where the HC component of the pressure dominates its electrostatic component, the SCDH-I and CCDH predictions roughly equivalent to the second order virial expansion of the CS pressure Buyuk2024 both agree with MC data up to the packing fraction $\eta\approx 0.1$. However, at the solute sizes $d<\ell_{\rm B}$ of Fig. 3(g)-(j) where the attractive energy component of the pressure embodying charge correlations becomes comparable with its repulsive HC component, the SCDH-I approach provides a more accurate prediction than the CCDH theory. In particular, at the substantially small solute sizes of Figs. 3(i)-(j) characterized by strong electrostatic coupling conditions, the SCDH-I formalism can reproduce with reasonable accuracy the plateau of the osmotic coefficient driven by the pronounced competition between opposite charge attraction and core repulsion over a broad concentration range.
III.2 Comparison of different splitting schemes
@@
The systematic comparison of the present formalism with simulations involving different salt valencies, concentrations, and HC sizes showed that the splitting of the short- and long-range interactions enables the SCDH-I formalism to handle electrostatic correlations in a significantly more accurate fashion than the CCDH approach. Here, we compare the underlying splitting scheme associated with the filter operator of fourth differential order (15) with the approach involving the inverse kernel $v^{-1}_{\rm l}(\mathbf{r},\mathbf{r}^{\prime})=\left(1-\sigma^{2}\nabla^{2}\right)v_{\rm c}^{-1}(\mathbf{r},\mathbf{r}^{\prime})$ including the filter operator of second differential order employed by Santangelo Santangelo , and the splitting approach of Ref. sp using error functions. Together with the constraint (10), these two different choices yield respectively the following short-range potentials,
@@
In Figs. 4(a)-(c), we show that the splitting lengths $\sigma$ obtained from the solution of Eq. (79) with these three distinct splitting potentials are significantly close, and they exhibit qualitatively similar behavior with the variation of the packing fraction or equivalently the salt concentration. Namely, in the HC- (electrostatically) dominated regimes of large (small) solute sizes, salt increment reduces (raises) the splitting length $\sigma$ towards the limit value of $\sigma\sim d/2$. Fig. 4(c) also shows that below a characteristic packing fraction, the splitting length $\sigma$ vanishes, implying that the SCDH-I formalism converges to the CCDH approach. Most importantly, in Fig. 4(d), one notes that the excess energy curves computed with the three different splitting potentials are considerably close. This indicates that the specific choice of the splitting scheme does not play a substantial role in the calculation of the bulk thermodynamic functions.
@@
III.3 Overscreening and underscreening
We evaluate here the accuracy of our formalism in predicting the deviation of the macromolecular interaction range from the DH length exp . Due to the high order of the moment integral $I^{(4)}_{i}$ in Eq. (58), in the calculation of the dressed charge, the uncertainties originating from the numerical inversion of the Fourier-transformed Green’s function (74) induce substantially large errors. In order to avoid this complication, we evaluated the screening parameter (62) within the quadratic splitting scheme of Eq. (86) enabling the analytical inversion of Eq. (74) as

### v2

@@ Abstract
ABSTRACT: We predict the thermodynamic behavior of bulk electrolytes from an ionic Hard-Core (HC) size-augmented self-consistent formalism incorporating asymmetrically the short- and long-range ion interactions via their virial and cumulant treatment, respectively. The characteristic splitting length separating these two ranges is obtained from a variational equation solved together with the Schwinger-Dyson (SD) equations. Via comparison with simulation results from the literature, we show that the asymmetric treatment of the distinct interaction ranges significantly extends the validity regime of our previously developed purely cumulant-level Debye-Hückel (DH) theory. Namely, for monovalent solutions with typical ion sizes, the present formalism can accurately predict up to molar concentrations the liquid pressure dominated by HC interactions, the internal energies driven by charge correlations, and the local ion distributions governed by the competition between HC and electrostatic interactions. We evaluate as well the screening length of the liquid and investigate the deviations of the macromolecular interaction range from the DH length. In fair agreement with simulations and experiments, our theory is shown to reproduce the overscreening and underscreening effects occurring respectively in submolar mono- and multivalent electrolytes.
@@ I Introduction
In order to extend the electrostatic coupling regime covered by the CCDH formalism, in this article, we develop a self-consistent DH (SCDH) theory of bulk electrolytes based on the asymmetric treatment of the short- and long-range electrostatic fluctuations. In Sec. II, we derive the field-theoretic partition function of the liquid embodying the corresponding splitting together with the pairwise HC interactions, and obtain a new variational equation solved by the splitting parameter separating these distinct interaction ranges. Then, we derive the SD equations relating this variational identity to the physical parameters of the system. Finally, we calculate the total correlation function required for the computation of the thermodynamic functions and the screening length of the electrolyte. The thermodynamic averages involved in these formally exact identities are evaluated via an hybrid calculation scheme that consists in incorporating asymmetrically the electrostatic fluctuations of short- and long-wavelengths within the virial and cumulant approximations, respectively. This scheme corresponding to a variational upgrade of our electrostatic WC-level CCDH formalism Buyuk2024 allows to avoid the cumulant-level WC treatment of the short-range ion interactions.
In Sec. III, we compare our predictions with hypernetted-chain (HNC) and MC data for various ion sizes, valencies, and concentrations. We find that the asymmetric variational treatment of the distinct interaction modes significantly improves the accuracy of the CCDH formalism in predicting the HC-dominated osmotic pressures, the electrostatic correlation-driven internal energies, and the ion pair distributions governed by the competing HC and electrostatic interactions. This upgrade is the main progress of our work. Via additional comparison with simulations and experiments, we equally show that our approach can also reproduce with reasonable accuracy the underscreening and overscreening effects occurring in mono- and multivalent solutions.
@@ II.6.1 Cumulant-level calculation of the ion fugacities
$-t\Lambda_{i}q_{i}^{2}\int\frac{\mathrm{d}^{3}\mathbf{r}_{1}\mathrm{d}^{3}\mathbf{r}_{2}}{2}\left[G_{{\rm l},0}^{-1}-v_{\rm{{\color[rgb]{0,0,0}l}}}^{-1}\right]_{\mathbf{r}_{1},\mathbf{r}_{2}}$
@@
with the rescaled fugacity $\Lambda_{i}=\lambda_{i}\;e^{-q_{i}^{2}\delta G_{{\rm l},0}(0)/2}$ and the Mayer function defined as
@@
$+tq_{i}^{2}n_{i}\int\frac{\mathrm{d}^{3}\mathbf{r}_{1}\mathrm{d}^{3}\mathbf{r}_{2}}{2}\left[G_{{\rm l},0}^{-1}-v_{\rm{{\color[rgb]{0,0,0}l}}}^{-1}\right]_{\mathbf{r}_{1},\mathbf{r}_{2}}$
@@ III Results and discussion
In this section, we compare the predictions of the present formalism with numerical and experimental data from the literature. For the interpretation of the results reported herein, it is essential to note that the CCDH theory of Ref. Buyuk2024 follows from the present formalism in the limit of vanishing splitting parameter $\sigma\to 0$ where the short-range potential (18) in Eq. (II.6.2) vanishes ($v_{\rm s}(r)\to 0$), and the Gaussian component of the long-range correlator reduces to the electrostatic WC-level DH potential, i.e. $G_{{\rm l},0}(r)\to v_{\rm DH}(r)=\ell_{\rm B}e^{-\kappa_{0}r}/r$, where $\kappa_{0}=\sqrt{8\pi\ell_{\rm B}n_{i}}$ is the DH screening parameter. Thus, the CCDH theory corresponds to the electrostatic WC limit of the SCDH approach. This implies that the pure HC interactions are embodied by the SCDH formalism at the same level of accuracy as the CCDH theory.
III.1 Radial distribution and charge density functions
Fig. 1 compares the theoretical predictions for the pair distribution functions around a central monovalent cation with MC simulation results (red symbols) at the ion size $d\approx 2.38$ and concentration $n_{i}\approx 1.17$ M. One sees that the cation-cation distribution function underestimated by the DH approximation $g_{ij}(r)\approx 1-q_{i}q_{j}v_{\rm DH}(r)$ (dotted curve) is accurately reproduced by both the CCDH (dashed curve) and the SCDH formalism (solid curve). In the case of the cation-anion distribution function, while the CCDH approach agrees with MC data only qualitatively, the simulation result is reproduced by the SCDH formalism with quantitative accuracy.
Figure 1: (Color online) Pair distribution functions associated with a central monovalent cation at the packing fraction $\eta=\pi n_{i}d^{3}/3=0.01$ and the reduced ion size $d/\ell_{\rm B}=1/3$. At the temperature $T=298$ K and the dielectric constant $\varepsilon_{\rm w}=78.5$, these parameters correspond to the ion size $d\approx 2.38$ and density $n_{i}\approx 1.17$ M. The MC data are from Fig. 1 of Ref. NetzMC .
In Figs. 2(a)-(c), the theoretical charge densities associated with a central cation are compared with HNC predictions at the larger ion size $d=4.6$ characterized by substantial HC correlations. One sees that the DH prediction ignoring the HC ion size is qualitatively inaccurate in all three plots. Then, the CCDH approach agrees qualitatively with the HNC data at the dilute concentration $n_{i}=0.1$ M, but significantly deviates from the HNC result at larger concentrations. Finally, the SCDH formalism agrees quantitatively with the HNC data in the entire concentration regime covered by the figure.
Figure 2: (Color online) Charge density $\rho_{\rm c}(r)=q_{+}n_{+}H_{i+}(r)+q_{-}n_{-}H_{i-}(r)$ around a central (a)-(c) cation ($i=+$) and (d)-(f) anion ($i=-$) in 1:1 liquids. In (a)-(c), the HNC data are from Fig. 6 of Ref. Kj2020 . The temperature and the dielectric permittivity are $T=298$ K and $\varepsilon=78.4$. The remaining parameters are provided in the legends. In (d)-(f), the MC data are from Fig. 8.B of Ref. NetzMC . From top to bottom, the packing fractions are $\eta=0.01$, $0.03$, and $0.15$. The reduced ion size is $d/\ell_{\rm B}=2/9$. The corresponding physical parameters are provided for $T=298$ K and $\varepsilon=78.5$ in the legends.
Figs. 2(d)-(f) display the charge distributions around a central anion at the significantly smaller HC size $d=1.59$ and large molar concentrations. This regime is characterized by the competition between pronounced HC interactions and substantial electrostatic correlations induced by the close interionic approach. One sees that the SCDH formalism improves the accuracy of the DH and CCDH predictions in a sizable fashion, and agrees quantitatively with the MC data over the entire concentration range considered in the figure. In particular, at the atypically large ion concentration of Fig. 2(f) characterized by a weak charge inversion (CI) effect, the SCDH formalism corrects the CCDH prediction by an order of magnitude, and reproduces the CI with reasonable accuracy.
@@ III.1 Radial distribution and charge density functions
Figure 3: (Color online) (a)-(c) Internal energy (83) and (b)-(d) pressure (84) of 1:1 solutions against salt concentration at two different HC sizes indicated in the legends. The temperature and the dielectric constant of the liquid are $T=300$ K and $\varepsilon_{\rm w}=78.5$. Solid curves: SCDH theory. Dashed curves: CCDH formalism of Ref. Buyuk2024 . Dotted curves: DHLLs. Red squares: HNC data from Figs. 3 and 4 of Ref. AttardPRE .
III.2 Excess energy and pressure
@@
The energy density and pressure of an electrolyte mixture can be expressed in terms of the total correlation function (80) considered in the previous part as Buyuk2024
@@
Figs. 3(a)-(d) compare the excess energy (83) and pressure (84) of the SCDH approach with HNC data for 1:1 liquids at two different ion sizes. The plots also display the CCDH predictions, and the DH limiting laws (DHLLs),
$\beta E_{{\rm ex,DH}}=-\frac{\kappa_{0}^{3}}{8\pi};\hskip 14.22636pt\beta P_{\rm DH}=\sum_{i=1}^{p}n_{i}-\frac{\kappa_{0}^{3}}{24\pi}.$ (85)
@@
In consistency with the ion distribution plots in Figs. 1-2, these energy and pressure plots show that the variational scheme at the basis of the SCDH formalism significantly improves the agreement of the CCDH theory with HNC data. Namely, while the CCDH formalism extends the validity of the DHLLs from $n_{i}\sim 10^{-2}$ M to $n_{i}\sim 0.2$ M, the SCDH results agree quantitatively with HNC data up to $n_{i}\sim 1.0$ M. Hence, for monovalent ions with typical hydration radii $d\sim 4$, the asymmetric treatment of the short- and long-range ion interactions enables the SCDH approach to accurately predict both the charge correlation-driven internal energy and the HC-dominated liquid pressure up to the molar concentration regime.
@@
Fig. 4(a) confronts the theoretical internal energy predictions for 1:1 solutions with MC simulations over a larger concentration regime. The quantitative accuracy of the SCDH approach up to $n_{i}\sim 1.0$ M is equally confirmed by this plot. At larger concentrations $n_{i}\gtrsim 2.0$ M where the SCDH prediction overestimates the MC data, the qualitative agreement of the present formalism with simulations is significantly better than that of the CCDH approach exhibiting a trend reversal, and an unphysical rise of the energy previously reported in Ref. Buyuk2024 .
@@
Figs. 4(b)-(d) show that for multivalent electrolytes characterized by stronger electrostatic coupling, the quantitative agreement of the SCDH formalism with MC data is limited to submolar concentrations. However, in the vicinity of the molar concentration regime, the SCDH approach can qualitatively reproduce the trend and magnitude of the simulation data with a considerably better accuracy than the CCDH formalism.
Figure 4: (Color online) Internal energy (83) against salt concentration. The ion valencies and sizes are indicated in the legends. The temperature and the dielectric constant are $T=298$ K and $\varepsilon_{\rm w}=78.5$. The curves have the same signification as in Fig. 3. The MC data (red squares) are from (a) Table I, (b) II, (c) III, and (d) IV of Ref. Valleau .
Figure 5: (Color online) (a)-(e) Internal energy (83) and (f)-(j) pressure (84) of monovalent ions against their packing fraction (lower axis) and concentration (upper axis) at the temperature $T=298$ K, the dielectric permittivity $\varepsilon_{\rm w}=78.5$, and the HC sizes indicated in the legends rem1 . MC data are from (a)-(e) Table I and (f)-(j) Table II of Ref. NetzMC .
@@
We investigate now the effect of HC size on the accuracy of the SCDH formalism. In Fig. 5(a), the comparison of the energy curves with MC data shows that at the particularly large solute size $d=14.29\;{\rm{\AA}}>\ell_{\rm B}\approx 7.1$ where the strong attenuation of opposite charge attraction by HC repulsion gives rise to weak electrostatic coupling conditions, the SCDH formalism brings a moderate correction to the CCDH approach. Then, in the typical case of smaller ion sizes $d<\ell_{\rm B}$ of Figs. 5(b)-(e) where the shorter interionic approach distances enhance the strength of the electrostatic correlations, the SCDH formalism substantially improves the quantitative accuracy of the CCDH approach, extending its validity limit by several factors into the molar concentration regime.
@@
In Figs. 5(f)-(j), we illustrate the effect of solute size on the liquid pressure (84). At the ion size $d=14.29$ where the HC component of the pressure dominates its electrostatic component, the SCDH and CCDH predictions roughly equivalent to the second order virial expansion of the CS pressure Buyuk2024 both agree with MC data up to the packing fraction $\eta\approx 0.1$. However, at the solute sizes $d<\ell_{\rm B}$ of Fig. 5(g)-(j) where the attractive energy component of the pressure embodying charge correlations becomes comparable with its repulsive HC component, the SCDH approach provides a more accurate prediction than the CCDH theory. In particular, at the substantially small solute size of Fig. 5(i) characterized by strong electrostatic coupling conditions, the SCDH formalism can reproduce with reasonable accuracy the plateau of the osmotic coefficient driven by the pronounced competition between opposite charge attraction and core repulsion over a broad concentration range.
III.3 Comparison of different splitting schemes
@@
The systematic comparison of the present formalism with simulations involving different salt valencies, concentrations, and HC sizes showed that the splitting of the short- and long-range interactions enables the SCDH formalism to handle electrostatic correlations in a significantly more accurate fashion than the CCDH approach. Here, we compare the underlying splitting scheme associated with the filter operator of fourth differential order (15) with the approach involving the inverse kernel $v^{-1}_{\rm l}(\mathbf{r},\mathbf{r}^{\prime})=\left(1-\sigma^{2}\nabla^{2}\right)v_{\rm c}^{-1}(\mathbf{r},\mathbf{r}^{\prime})$ including the filter operator of second differential order employed by Santangelo Santangelo , and the splitting approach of Ref. sp using error functions. Together with the constraint (10), these two different choices yield respectively the following short-range potentials,
@@
In Figs. 6(a)-(c), we show that the splitting lengths $\sigma$ obtained from the solution of Eq. (79) with these three distinct splitting potentials are significantly close, and they exhibit qualitatively similar behavior with the variation of the packing fraction or equivalently the salt concentration. Namely, in the HC- (electrostatically) dominated regimes of large (small) solute sizes, salt increment reduces (raises) the splitting length $\sigma$ towards the limit value of $\sigma\sim d/2$. Fig. 6(c) also shows that below a characteristic packing fraction, the splitting length $\sigma$ vanishes, implying that the SCDH formalism converges to the CCDH approach. Most importantly, in Fig. 6(d), one notes that the excess energy curves computed with the three different splitting potentials are considerably close. This indicates that the specific choice of the splitting scheme does not play a substantial role in the calculation of the bulk thermodynamic functions.
@@
III.4 Overscreening and underscreening
We evaluate here the accuracy of our formalism in predicting the deviation of the macromolecular interaction range from the DH length exp ; safran ; andelman . Due to the high order of the moment integral $I^{(4)}_{i}$ in Eq. (58), in the calculation of the dressed charge, the uncertainties originating from the numerical inversion of the Fourier-transformed Green’s function (74) induce substantially large errors. In order to avoid this complication, we evaluated the screening parameter (62) within the quadratic splitting scheme of Eq. (86) enabling the analytical inversion of Eq. (74) as